OpenStudy (anonymous):

Write the recursive formula for the geometric sequence. a1 = -2 a2 = 8 a3 = -32

OpenStudy (anonymous):

\[a_1=-2,a_2=-4a_1=-4a_2\therefore a_{n+1}=-4a_n\]

OpenStudy (anonymous):

thats not my choices

OpenStudy (anonymous):

What are your choices?

OpenStudy (radar):

\[\sqrt{-2*8}=\sqrt{-16}=4i\] That seems to be the ratio?

OpenStudy (anonymous):

A. an = -4 + an-1 B. an = -2 + an-1 C. an = -2 • an-1 D. an = -4 • an-1

OpenStudy (phi):

when you write down a recursion, you can say a(n+1)= -4 a(n) or a(n)= -4 a(n-1) both mean the same thing: the next a term is -4 times the current a term. or the current a term is -4 times the previous a term so badrefs answer is correct.

OpenStudy (anonymous):

I don't think a recursion formula is dependent on the first term, is it? We can write the general formula \(a_n=-2{\left(-4\right)}^n\).

OpenStudy (radar):

Thanks phi for straighten that out. I thought the a1,a2 were the label for the sequence members, didn't even realize it was part of the sequence terms.

OpenStudy (anonymous):

Write the explicit formula for the geometric sequence. a1 = -5 a2 = 20 a3 = -80

OpenStudy (phi):

@badreferences if you believe wikipedia http://openstudy.com/study#/updates/4f687d3ee4b0f81dfbb562ac In mathematics, a recurrence relation is an equation that recursively defines a sequence, once one or more initial terms are given: each further term of the sequence is defined as a function of the preceding terms.

OpenStudy (phi):

Write the explicit formula for the geometric sequence. a1 = -5 a2 = 20 a3 = -80 Here they want the closed form: a(n)= -5*(-4)^(n-1)

OpenStudy (anonymous):

No, the recurrence relation doesn't necessarily have to include the previous term according to Wikipedia http://en.wikipedia.org/wiki/Logistic_map . I think it's fair to say all the solutions have been provided, then.

OpenStudy (phi):

In this problem, they are making a distinction between a recurrence and a closed form.